How can I use the vector product to find the angle between the vectors:
v= 6i - 3j + k and w= 2i + 2j - k. ???
Any help would be great.
-Nick

Answer by Alan3354(69443)

How can I use the vector product to find the angle between the vectors:
v= 6i - 3j + k and w= 2i + 2j - k.
------------
The dot product is the product of the 2 magnitudes * cos of the angle between the vectors.
|v| = sqrt(36 + 9 + 1) = sqrt(46)
|w| = sqrt(4 + 4 + 1) = 3
----------------
v dot w = 6*2 + -3*2 + -1 = 5
5 = 3*sqrt(46)*cos(A)
cos(A) = 5/(3sqrt(46)) =~ 0.2457366
Angle =~ 75.7746 degs